the sum of three integers is 308. the first and second exceed the third by 84. the third integer is 10 less than the first. find the three integers.
Answer by savvyhush23(50)

the sum of three integers is 308. the first and second exceed the third by 84. the third integer is 10 less than the first. find the three integers.
.
.
Let x,y and z be the intergers,
.
.
sum of three integers is 308: x + y + z = 308 equation (1)
the first and second exceed the third by 84: x + y = z + 84 equation (2)
third integer is 10 less than the first: z = x - 10 equation (3)
.
substitute (2) to (1): (z + 84) + z = 308 ; 2z = 224
substitute (3): 2(x - 10) = 224 ; x = 122
Therefore, x = 122, y = 74 , z = 112